OpenStudy (anonymous):

Find an equation of the lines passing thru each pair of points. Write eqution in form Ax+Bx=C... Here are the points (2,3) and (-1,-1)

OpenStudy (anonymous):

Find slope between two points Use this (y-y1)=m(x-x1) plug in slope in m

OpenStudy (anonymous):

You can rearrange terms to get it into standard form

OpenStudy (anonymous):

I get -2x+y=-1, but in the back of the book its 4x-3y=-1...HOW